Grade 11Algebra

if there is a common root of x2+px+q=0and x2+qx+pr=0,then proof that p+q+r=0

Debasish Bhattacharya
7 Years agoGrade 11

1 Answer

Samyak Jain
7 Years ago
The equations given are slightly wrong. By solving them we get wrong result.
Equations should be x2 + px + qr = 0 and x2 + qx + pr = 0.
Concept : Let ax2 + b1 x + c1 = 0 and ax2 + b2 x + c2 = 0.
Let \alpha be the common root of the above equations.
By cross multiplication method, we get
\alpha2 / (bc2 – bc1)  = \alpha / (c1 a2 – c2 a1)  =  1 / (a1b2 – a2 b1)
\therefore  \alpha = (bc2 – bc1) / (c1 a2 – c2 a1)  =  (c1 a2 – c2 a1) / (a1b2 – a2 b1)      ...(1)
Consider (bc2 – bc1) / (c1 a2 – c2 a1)  =  (c1 a2 – c2 a1) / (a1b2 – a2 b1
\Rightarrow (bc2 – bc1)(a1b2 – a2 b1) = (c1 a2 – c2 a1)2  …(2), which is the condition for common root.
Comparing given equations x2 + px + qr = 0 and x2 + qx + pr = 0 with 
  ax2 + b1 x + c1 = 0 and ax2 + b2 x + c2 = 0,
we get  a= 1,  b1 = p,  c1 = qr,  a2 = 1,  b2 = q,  c2 = pr.
Putting these values in (2), we get
(p2r – q2r)(q – p) = (qr – pr)2  i.e.  r(p2 – q2)(q – p) = r2(q – p)2 ,now r(q – p) will get cancelled from both sides
(p2 – q2) = r(q – p)  or  (p + q)(p – q) = r(q – p)
\therefore (p + q) = – r  \Rightarrow  p + q + r = 0.
(Note : You can memorise result (2) to solve these type of questions.)